xxxmariahlulbaddiexx:

Write a linear equation that would be parallel to f(x) = -3x +4

snowflake0531:

Change the y-intercept and it'll be parallel for example if you had y=-2x+3, y=-2x+5 would be parallel

Extrinix:

@snowflake0531 wrote:
Change the y-intercept and it'll be parallel for example if you had y=-2x+3, y=-2x+5 would be parallel
She's basically saying to take the b variable in y=mx+b and change it slight, so examples for your equation would be, f(x)=-3x+6 f(x)=-3x-7 f(x)=-3x+\(\dfrac{2}{3}\) etc.